Consider a Keynesian model: 

Full employment output = R100 million 

Tax rate = 0,25 

Investment = R40 million 

Autonomous consumption = R30 million 

Marginal propensity to consume = 0,8 

3.28 The value of the multiplier is … 

[1] 2 

[2] 1.67 

[3] 2.5 

[4] 4 


3.29 The equilibrium level of income is … 

[1] R70 million. 

[2] R175 million. 

[3] R280 million. 

[4] R140 million. 

3.30 To bring about full employment, government spending should be … 

[1] -R30 million. 

[2] -R72 million. 

[3] R30 million. 

[4] R75 million. 


1
Expert's answer
2021-07-15T12:27:04-0400

a) The value of the multiplier is [4.] 4.

Multiplier = "\frac{1}{(1 - MPC)}"

MPC = 0.8

Multiplier = "\frac{1}{(1 - 0.8)} = \frac{1}{0.2} = 5"

The value of the multiplier = 5

 

b.). The equilibrium level of income is [2.] R175 million.

At equilibrium: AD = AS

Y = C + I + G

C = 30 + 0.8YD = 30 + 0.8(Y – T) = 30 + 0.8(Y – 0.25Y)

C = 30 + 0.8(Y – 0.25Y)

I = 40

Y = 30 + 0.8(Y – 0.25Y) + 40

Y = 30 + 0.6Y + 40

Y – 0.6Y = 30 + 40

0.4Y = 70

Y = 175

The equilibrium level of income is R175 million.

 

c.). To bring about full employment, government spending should be [1] -R30 million.

GDP needs to be decreased by (175 – 100 = 75 million) to bring about full employment.

Plug the full employment level of 100 million same as Y and solve for government spending.

Y = C + I + G

100 = 30 + 0.8(Y – 0.25Y) + 40 + G

100 = 30 + 0.8(100 – 0.25(100) + 40 + G

100 = 30 + 60 + 40 + G

100 = 130 + G

100 – 130 = G

G = -30

To bring about full employment, government spending should be -R30 million.